A car travels the lenght of the turnpike that is 600 km long in 6 hours and 50 minutes. what is the cars speed in km/h? and m/s?
Answer by checkley77(12569)

D=RT OR R=D/T
R=600/(6+5/6)
R=600/[(36+5)/6]
R=600/(41/6)
R=600*6/41
R=3,600/41
R=87.8 KPH. ANS.
PROOF:
600=87.8(6+5/6)
600=87.8*6.8333
600~600